OpenStudy (anonymous):

A Childish Question... But really wanna know the Steps b/w them. (5/9)[exponent]-3/2 ………………27/5√5 (1/2)[exponent]-3/2 ………………√8

OpenStudy (mertsj):

\[\frac{5}{9}^{\frac{-3}{2}}\]

OpenStudy (mertsj):

\[(\frac{9}{5})^{\frac{3}{2}}\]

OpenStudy (mertsj):

\[\frac{3^{3}}{\sqrt{5}^{3}}\]

OpenStudy (mertsj):

\[\frac{27}{\sqrt{5}\sqrt{5}\sqrt{5}}\]

OpenStudy (mertsj):

\[\frac{27}{5\sqrt{5}}\]

OpenStudy (mertsj):

\[(\frac{1}{2})^{\frac{-3}{2}}\]

OpenStudy (mertsj):

\[2^{\frac{3}{2}}\]

OpenStudy (mertsj):

\[\sqrt{2}\sqrt{2}\sqrt{2}=\sqrt{8}\]

OpenStudy (anonymous):

Got it.. Thank you very much :)

OpenStudy (mertsj):

yw